# Optimal Strategy Revision in Population Games:

### v1

@@ Abstract
This paper investigates the design of optimal strategy revision in Population Games (PG) by establishing its connection to finite-state Mean Field Games (MFG). Specifically, by linking Evolutionary Dynamics (ED) – which models agent decision-making in PG – to the MFG framework, we demonstrate that optimal strategy revision can be derived by solving the forward Fokker-Planck (FP) equation and the backward Hamilton-Jacobi (HJ) equation, both central components of the MFG framework. Furthermore, we show that the resulting optimal strategy revision satisfies two key properties: positive correlation and Nash stationarity, which are essential for ensuring convergence to the Nash equilibrium. This convergence is then rigorously analyzed and established. Additionally, we discuss how different design objectives for the optimal strategy revision can recover existing ED models previously reported in the PG literature. Numerical examples are provided to illustrate the effectiveness and improved convergence properties of the optimal strategy revision design.
Index Terms:
Population Games, Mean Field Games, Multi-Agent Strategic Interactions
@@ I Introduction
Evolutionary Dynamics (ED) are decision-making models that guide a large population of agents to learn strategies that form a Nash equilibrium in Population Games (PG). These models are characterized by strategy revision protocols, which define how agents revise between strategies based on the payoffs received from their current strategy choices. While various revision protocols have been studied, a key question remains unanswered: which protocol is the most effective for achieving the equilibrium, given a specified objective?
Recent advances in the PG literature have introduced higher-order dynamics models for refining payoff determinations [1, 2]. These models, for example, can be designed to filter out noise in perceived payoffs [3] or to mitigate the adverse effects of time delays on the convergence to the Nash equilibrium [4]. In this work, leveraging the Mean Field Games (MFG) framework, we propose a novel approach to designing a payoff dynamics model that optimally balances rapid convergence to the Nash equilibrium with the penalties incurred during strategy revision.
@@
Solutions for standard MFG problems are characterized by two equations: a backward Hamilton-Jacobi equation, which describes the evolution of the optimal cost-to-go, and a forward Fokker-Planck-Kolmogorov equation, which captures the distribution of agents’ states. Although both MFG and PG focus on decision-making in large populations, their primary difference lies in the structure of their strategy sets: MFG typically involves a continuous strategy set, as seen in differential games of the mean-field type [5], whereas PG is defined over a finite number of strategies [6].
@@
The study by [7] introduces a finite-state MFG framework, laying the groundwork for establishing a connection between MFG and PG. Building on this foundation, the primary contribution of our work is to formally establish the relationship between MFG and PG frameworks. Specifically, we demonstrate the connection between the Fokker-Planck (FP) equation in finite-state MFG and the ED model in PG. Leveraging this connection, we show that the optimal solution in the finite-state MFG framework corresponds to an optimal strategy revision protocol in PG, with payoff dynamics derived from the backward Hamilton-Jacobi (HJ) equation. To the best of our knowledge, this is the first work to formally establish a direct connection between MFG and ED models in PG.
@@
The contributions of this paper can be summarized as follows: $i)$ We establish a connection between the finite-state MFG introduced in [7] and the ED models presented in [6, 8]. This is achieved by introducing a dynamic model for payoff vectors analogous to the backward HJ equation in MFG. $ii)$ We demonstrate that, under the proposed payoff dynamics model, changes in the distribution of the agents’ strategy selections exhibit positive correlation with the payoff vector. This behavior aligns with the positive correlation property observed in conventional ED models and is critical for ensuring convergence to the Nash equilibrium. Additionally, we establish the link between the steady-state solution of the MFG model and the Nash equilibrium in PG. The convergence of the population state to the Nash equilibrium is rigorously discussed. $iii)$ We show that existing ED models, including Smith, replicator, and projection dynamics, can be derived as special cases of our approach.
This paper is organized as follows: §$\S$§II provides the problem statement and a brief overview of preliminaries on PG and finite-state MFG. §$\S$§III presents the main results of this paper. Illustrative examples are discussed in §$\S$§IV. Finally, conclusions are drawn in §$\S$§V.
@@ II-A Population Games and Problem Formulation
Consider a large population of decision-making agents, each selecting a strategy from the set $\mathbb{S}=\{1,\cdots,n\}$. At any time $t\in[t_{0},T]$ with $t_{0}<T$, agents may revise their strategy choices. The vector-valued function $x:[t_{0},T]\to\Delta$ represents the state of the population, describing the distribution of agents’ strategy selections at a given time $t$. The set $\Delta$, known as the probability simplex, is defined as $\Delta=\{x\in\mathbb{R}_{+}^{n}\,|\,\sum_{i\in\mathbb{S}}x_{i}=1\}$, representing the set of all feasible population states. The $i$-th entry $x_{i}(t)$ of $x(t)$ specifies the proportion of agents selecting strategy $i\in\mathbb{S}$ at time $t$.
@@
A population state $x^{*}\in\Delta$ is called a Nash equilibrium of $F$ if the following holds for all $i$ in $\mathbb{S}$:
@@
The set of all Nash equilibria is denoted by $\mathbb{NE}(F)$.
@@
A population game $F$ is called contractive if the following inequality holds for all $x,y$ in $\Delta$:
$\left(F(x)-F(y)\right)^{\top}\left(x-y\right)\leq 0.$ (1)
Agents are allowed to revise their strategy selections at the arrival times of i.i.d. Poisson processes [6]. Let $\rho_{ji}(p,x)$ denote the strategy revision protocol, which specifies the likelihood that a strategy-revising agent switches from strategy $j$ to strategy $i$. For instance, the Smith protocol is defined as $\rho_{ji}(p,x)=[p_{i}-p_{j}]_{+}$, where $[\cdot]_{+}=\max(0,\cdot)$. As shown in [6], under the assumptions of a homogeneous population – where all agents adopt the same revision protocol – and an infinitely large population, the population state $x(t)$ evolves according to the following ordinary differential equation, commonly referred to as the Evolutionary Dynamics (ED):
@@
where $p(t)=F(x(t))$. Let $V$ be the vector field defining $\dot{x}(t)=V(p(t),x(t))$, where $x(t)=(x_{1}(t),\cdots,x_{n}(t))$. As discussed in [9], the following key properties of ED are crucial in establishing convergence to the Nash equilibrium.
@@
ED satisfies $i)$ the Nash stationarity if $V(F(x),x)=0$ holds if and only if $x\in\mathbb{NE}(F)$, and $ii)$ the positive correlation if $V(p,x)\neq\mathbf{0}$ implies $p^{\top}V(p,x)>0$.
@@
$\dot{p}_{i}(t)=G_{i}(t,p(t),x(t)),~{}i\in\mathbb{S},$ (3)
that maximize the agents’ long-term payoffs. The problem is formulated as an optimal control problem for a reference agent starting with each strategy $i$. The reference agent is defined as a randomly selected agent whose strategy selection at time $t_{0}$ is $i$ [^1] In population games, agents are non-atomic, meaning their individual identities are negligible; only their strategy choices matter. Consequently, the finite time-horizon payoff function (4) describes the long-term payoffs received by agents who begin with strategy $i$ under the revision protocol $\rho$.
Main Problem: Consider an optimal control problem where a reference agent initially selects strategy $i$ at time $t_{0}$ and subsequently revises its strategy according to the strategy revision protocol. The objective is to design $\rho=(\rho_{ij})_{i,j\in\mathbb{S}}$ and $G=(G_{i})_{i\in\mathbb{S}}$ to maximize a finite time-horizon payoff function defined as: [^2] The expectation in (4) is taken with respect to the strategy selection $s(t)$, a random variable over $\mathbb{S}$ with the transition probability satisfying [10, §4]: $\lim_{\epsilon\to 0}\frac{1}{\epsilon}\mathbb{P}\left[s(t+\varepsilon)=j\,\big{|}\,s(t)=i\right]=\rho_{ij}(p(t),x(t))$.
$P_{i}(t_{0},\rho)=\mathbb{E}\Big{[}\textstyle\int_{t_{0}}^{T}\big{(}-\sum_{j\in\mathbb{S}}\frac{q_{s(t),j}(t)}{2}\rho_{s(t),j}^{2}(p(t),x(t))\\ +F_{s(t)}(x(t))\big{)}\mathrm{d}t+F_{s(T)}(x(T))\,\Big{|}\,s(t_{0})=i\Big{]},$ (4)
where $F=(F_{1},\cdots,F_{n})$ is the instantaneous payoff function of the underlying population game, $s(t)\in\mathbb{S}$ denotes the agent’s strategy at time $t$, and $x(t)$ and $p(t)$ are determined by (2) and (3), respectively.
The finite time-horizon payoff function $P_{i}(t_{0},\rho)$ accumulates payoffs over $[t_{0},T]$, including a penalty term proportional to the square of the protocol $\rho_{ji}$. The weight $q_{ij}(t),~{}i,j\in\mathbb{S}$ plays a crucial role in the optimal control problem. Reducing $q_{ij}(t)$ decreases the penalization for revising strategies from $i$ to $j$, thereby accelerating the dynamics in (2). The weight $q_{ij}(t)$ may depend on the population states, as discussed in §$\S$§III-A.
II-B Finite-State Mean-Field Games
Consider a large population of homogeneous agents interacting within a finite-state MFG framework over a time horizon $[t_{0},T]$ with $t_{0}<T$, as discussed in [7]. In this framework, the state of a reference agent is modeled by a continuous-time Markov chain defined over a finite state space $\mathbb{S}=\{1,\cdots,n\}$. State transitions are governed by an $n\times n$ transition matrix $\alpha(t)=(\alpha_{ij}(t))_{i,j\in\mathbb{S}}\in\mathbb{R}^{n\times n}$, satisfying the following conditions at each time $t$: $\alpha_{ij}(t)\geq 0,~{}\forall i,j\in\mathbb{S}$, $i\neq j$, and $\sum_{j\in\mathbb{S}}\alpha_{ij}(t)=0,~{}\forall i\in\mathbb{S}$. More formally, for $\varepsilon>0$, the transition probability is defined as
@@ II-B Finite-State Mean-Field Games
where $\alpha_{i}(t)=(\alpha_{i1}(t),\cdots,\alpha_{in}(t))$ is the $i$-th row of the transition matrix, $c:\mathbb{S}\times\Delta\times\mathbb{R}^{n}\to\mathbb{R}$ is the running cost, and $\psi:\mathbb{S}\times\Delta\to\mathbb{R}$ is the terminal cost.
For notational convenience, given $i$ in $\mathbb{S}$, let $\mathbb{R}_{i,0}^{n}$ denote the set of vectors $\mu=(\mu_{1},\cdots,\mu_{n})\in\mathbb{R}_{i,0}^{n}$ satisfying: $i)$ $\mu_{j}\geq 0$ if $j\neq i$, and $ii)$ $\sum_{j\in\mathbb{S}}\mu_{j}=0$. The optimal cost-to-go from time $t$ is defined as $w_{i}(t,x)=\min_{\alpha}J_{i}(t,\alpha)$.
The solution to the MFG problem is given by two coupled equations, as outlined in [7]. The first is the backward evolution of $w$, known as the Hamilton-Jacobi (HJ) equation:
@@
The second equation describes the forward evolution of $x(t)$, commonly referred to as the Fokker-Plank (FP) equation:
@@
where $(x_{0,1},\cdots,x_{0,n})\in\Delta$, and
$\alpha_{i}(t)=\textstyle\operatorname*{arg\,min}_{\mu\in\mathbb{R}_{i,0}^{n}}\Big{(}c(i,x(t),\mu)\\ +\sum_{j\in\mathbb{S}}\mu_{j}(w_{j}(t,x(t))-w_{i}(t,x(t)))\Big{)},$
representing the optimal state transition. For the existence of solutions to (7) and (9), see [7, Proposition 4].
The MFG system (7) and (9) may not always admit a stationary solution, defined as $\dot{w}(t,x(t))=\dot{x}(t)=0$, for instance, when $H_{i}>0$. In such cases, we define a stationary solution $(\bar{w},\bar{x})\in\mathbb{R}^{n}\times\Delta$ (modulo addition of a constant $\kappa$) as follows: For all $i$ in $\mathbb{S}$,
$H_{i}(\bar{w},\bar{x})$ $=\kappa$ (10a)
$\textstyle\sum_{j\in\mathbb{S}}\bar{\alpha}_{ji}\bar{x}_{j}$ $=0,$ (10b)
where $\bar{\alpha}_{i}=\operatorname*{arg\,min}_{\mu\in\mathbb{R}_{i,0}^{n}}(c(i,\bar{x},\mu)+\sum_{j\in\mathbb{S}}\mu_{j}(\bar{w}_{j}-\bar{w}_{i}))$.

### v2

@@ Abstract
This paper investigates the design of optimal strategy revision in Population Games (PG) by establishing its connection to finite-state Mean Field Games (MFG). Specifically, by linking Evolutionary Dynamics (ED)—which models agent decision-making in PG—to the MFG framework, we demonstrate that optimal strategy revision can be derived by solving the forward Fokker-Planck (FP) equation and the backward Hamilton-Jacobi (HJ) equation, both central components of the MFG framework. Furthermore, we show that the resulting optimal strategy revision, which maximizes each agent’s payoffs over a finite time horizon, satisfies two key properties: positive correlation and Nash stationarity, which are essential for ensuring convergence to the Nash equilibrium. This convergence is then rigorously analyzed and established. Additionally, we discuss how different design objectives for the optimal strategy revision can recover existing ED models previously reported in the PG literature. Numerical examples are provided to illustrate the effectiveness and improved convergence properties of the optimal strategy revision design.
@@ I Introduction
Large-population game frameworks are widely applicable across engineering domains such as multi-agent task allocation, demand response in smart grids, wireless communication networks, distributed control systems, and large-scale optimization problems [Park2019Payoff-Dynamic-, §III], [7823106], where numerous agents engage in strategic interactions. In particular, evolutionary dynamics (ED) describe decision-making models that enable large populations of agents to adapt and learn strategies that converge to a Nash equilibrium in population games (PG). Furthermore, mean field games (MFG) offer an analytically tractable framework for solving stochastic differential games involving large populations [Lasry2007, Caines].
Both PG and MFG provide rigorous frameworks for designing decision-making models applicable to scenarios involving finite but large populations of agents. Although they share this overarching objective, they differ fundamentally in their modeling perspectives. The PG framework focuses on deriving decision models for large populations of agents engaged in myopic decision-making based on instantaneous payoffs [Sandholm2010-SANPGA-2]. These models are governed by strategy revision protocols, which specify how agents switch between strategies based on the payoffs associated with their current strategy choices. In contrast, the MFG framework aims to compute solutions that optimize long-term cost functions dependent on the evolving population’s state in either deterministic or stochastic differential game settings, considered in the population limit [Lacker2016].
A substantial body of literature on PG has explored a variety of decision-making models and developed techniques for establishing their stability (e.g., [HOFBAUER20091665, HOFBAUER200747, Fox2013Population-Game, 9219202, 10933512, EDM_water_passivity]). Building on these foundations, recent advances have introduced higher-order dynamics models to refine payoff evaluations [Fox2013Population-Game, park2018cdc]. Such models can, for example, be designed to filter noise in perceived payoffs [Park2019Payoff-Dynamic-] or mitigate the adverse effects of time delays on convergence to the Nash equilibrium [Park2023LearningWD]. Other works, such as [9022871], have established the stability of certain classes of decision-making models in PG by uncovering connections between PG and finite game formulations and leveraging stability analysis techniques developed in the latter context.
Despite these significant advances, a fundamental question remains unresolved: which strategy revision protocol is most effective for achieving equilibrium under a given objective? In this work, we address this question by leveraging the MFG framework to propose a novel approach for designing payoff dynamics models that optimally balance rapid convergence to the Nash equilibrium with the penalties associated with strategy revisions. To the best of our knowledge, this is the first study to formally establish a direct connection between MFG and ED models in PG, thereby enabling the computation of optimal strategy revision protocols.
@@
Solutions for standard MFG problems are characterized by two equations: a backward Hamilton-Jacobi equation, which describes the evolution of the optimal cost-to-go, and a forward Fokker-Planck-Kolmogorov equation, which captures the distribution of agents’ states. Traditionally, although both MFG and PG focus on decision-making in large populations, their primary difference lies in the structure of their strategy sets: MFG typically involves a continuous strategy set, as seen in differential games of the mean-field type [CRC_Book_BaTe], whereas PG is defined over a finite number of strategies [Sandholm2010-SANPGA-2].
@@
The study by [Diogo_2013] introduces a finite-state MFG framework, laying the groundwork for establishing a connection between MFG and PG. Building on this foundation, the primary contribution of our work is to formally establish the relationship between MFG and PG frameworks. Specifically, we demonstrate the connection between the Fokker-Planck (FP) equation in finite-state MFG and the ED model in PG. Leveraging this connection, we show that the optimal solution in the finite-state MFG framework corresponds to an optimal strategy revision protocol in PG, with payoff dynamics derived from the backward Hamilton-Jacobi (HJ) equation.
@@
The contributions of this paper can be summarized as follows: $i)$ We establish a connection between the finite-state MFG introduced in [Diogo_2013] and the ED models presented in [Sandholm2010-SANPGA-2, SANDHOLM2015703]. This is achieved by introducing a dynamic model for payoff vectors analogous to the backward HJ equation in MFG. ii)ii) We demonstrate that, under the proposed payoff dynamics model, changes in the distribution of the agents’ strategy selections exhibit positive correlation with the payoff vector. This behavior aligns with the positive correlation property observed in conventional ED models and is critical for ensuring convergence to the Nash equilibrium. Additionally, we establish the link between the steady-state solution of the MFG model and the Nash equilibrium in PG. The convergence of the population state to the Nash equilibrium is rigorously discussed. $iii)$ We show that existing ED models, including Smith, replicator, and projection dynamics, can be derived as special cases of our approach.
[…]
@@ II-A Population Games and Problem Formulation
Consider a large population of decision-making agents, each selecting a strategy from the set $\mathbb{S}=\{1,\cdots,n\}$. At any time $t\in[t_{0},T]$ with $0\leq t_{0}<T$, agents may revise their strategy choices. The vector-valued function $x:[t_{0},T]\to\Delta$ represents the state of the population, describing the distribution of agents’ strategy selections at a given time $t$. The $i$-th entry $x_{i}(t)$ of $x(t)$ specifies the proportion of agents selecting strategy $i\in\mathbb{S}$ at time $t$. The set $\Delta$, known as the probability simplex, is defined as $\Delta=\{x\in\mathbb{R}_{+}^{n}\,|\,\sum_{i\in\mathbb{S}}x_{i}=1\}$, representing the set of all feasible population states. The relative interior of the simplex is given by $\mathrm{int}(\Delta)=\{x\in\Delta\,|\,x_{i}>0,\forall i\in\mathbb{S}\}$.
@@
A population state $x^{*}\in\Delta$ is called a Nash equilibrium of $F$ if the following holds for all $i\in\mathbb{S}$:
@@
The set of all Nash equilibria is denoted by $\mathbb{NE}(F)$. $\square$
@@
A population game $F$ is called contractive if the following inequality:
$\left(F(x)-F(y)\right)^{\top}\left(x-y\right)\leq 0$ (1)
holds for all $x,y\in\Delta$. $\square$
Agents are allowed to revise their strategy selections at the arrival times of i.i.d. Poisson processes [Sandholm2010-SANPGA-2]. Let $\rho_{ji}(p,x)$ denote the strategy revision protocol, which specifies the rate at which a strategy-revising agent switches from strategy $j$ to strategy $i$. For instance, the Smith protocol is defined as $\rho_{ji}(p,x)=[p_{i}-p_{j}]_{+}$, where $\cdot]_{+}=\max(0,\cdot)$. As shown in [Sandholm2010-SANPGA-2], under the assumptions of a homogeneous population—where all agents adopt the same revision protocol—and an infinitely large population, the population state $x(t)$ evolves according to the following ordinary differential equation, commonly referred to as the Evolutionary Dynamics (ED):
@@
where $p(t)=F(x(t))$. Let $V$ be the vector field defining $\dot{x}(t)=V(p(t),x(t))$, where $x(t)=(x_{1}(t),\cdots,x_{n}(t))$. As discussed in [HOFBAUER20091665], the following key properties of ED are crucial in establishing convergence to the Nash equilibrium.
@@
ED (2) satisfies $i)$ the Nash stationarity if $V(F(x),x)=0$ holds if and only if $x\in\mathbb{NE}(F)$, and ii)ii) the positive correlation if $V(p,x)\neq 0$ implies $p^{\top}V(p,x)>0$. $\square$
@@
$\dot{p}_{i}(t)=G_{i}(p(t),x(t)),~i\in\mathbb{S},$ (3)
that maximize the agents’ long-term payoffs. The problem is formulated as an optimal control problem for a reference agent initialized with strategy $i$. The reference agent is defined as a randomly selected individual whose strategy at the initial time $t_{0}$ is $i$. As in the standard framework of population games, each agent revises its strategy based on its own payoff. The effect of an agent’s strategic decisions on its future payoffs is captured through the population state $x(t)$ in the formulation. Consequently, the finite time-horizon payoff function is defined exclusively in terms of the current and anticipated future payoffs of the agent.
Main Problem: Consider an optimal control problem where a reference agent initially selects strategy $i$ at time $t_{0}$ and subsequently revises its strategy according to the strategy revision protocol. The objective is to design $\rho=(\rho_{ij})_{i,j\in\mathbb{S}}$ and $G=(G_{i})_{i\in\mathbb{S}}$ to maximize a finite time-horizon payoff function defined as: [^1] In population games, agents are non-atomic, meaning their individual identities are negligible; only their strategy choices matter. Consequently, the finite time-horizon payoff function (4) describes the long-term payoffs received by agents who begin with strategy $i$ under the revision protocol $\rho$.
$P_{i}(t_{0},\rho)=\mathbb{E}\Big[\textstyle\int_{t_{0}}^{T}\big(-\sum_{j\in\mathbb{S}}\frac{q_{s(t),j}(x(t))}{2}\rho_{s(t),j}^{2}(p(t),x(t))\\ +F_{s(t)}(x(t))\big)\mathrm{d}t+F_{s(T)}(x(T))\,\Big|\,s(t_{0})=i\Big],$ (4)
where $F=(F_{1},\cdots,F_{n})$ is the instantaneous payoff function of the underlying population game, $s(t)\in\mathbb{S}$ denotes the agent’s strategy at time $t$, and $x(t)$ and $p(t)$ are determined by (2) and (3), respectively. We assume that $F$ is continuously differentiable on $\Delta$. The expectation in (4) is taken with respect to the strategy selection $s(t)$, a random variable over $\mathbb{S}$ with the transition probability satisfying [park2018cdc, §4]: $\lim_{\epsilon\to 0}\frac{1}{\epsilon}\mathbb{P}\left[s(t+\varepsilon)=j\,\big|\,s(t)=i\right]=\rho_{ij}(p(t),x(t))$.
[…]
II-B Finite-State Mean Field Games
Consider a large population of homogeneous agents interacting within a finite-state MFG framework over a time horizon $t_{0},T]$ with $0\leq t_{0}<T$, as discussed in [Diogo_2013]. In this framework, the state of a reference agent is modeled by a continuous-time Markov chain defined over a finite state space $\mathbb{S}=\{1,\cdots,n\}$. State transitions are governed by an $n\times n$ transition matrix $\alpha(t)=(\alpha_{ij}(t))_{i,j\in\mathbb{S}}\in\mathbb{R}^{n\times n}$, satisfying the following conditions at each time $t$: $\alpha_{ij}(t)\geq 0,~\forall i,j\in\mathbb{S}$, $i\neq j$, and $\sum_{j\in\mathbb{S}}\alpha_{ij}(t)=0,~\forall i\in\mathbb{S}$. More formally, for $\varepsilon>0$, the transition probability is defined as
@@ II-B Finite-State Mean Field Games
where $\alpha_{i}(t)=(\alpha_{i1}(t),\cdots,\alpha_{in}(t))$ is the $i$-th row of the transition matrix, $c:\mathbb{S}\times\Delta\times\mathbb{R}^{n}\to\mathbb{R}$ is the running cost, and $\psi:\mathbb{S}\times\Delta\to\mathbb{R}$ is the terminal cost. For notational convenience, given $i\in\mathbb{S}$, let $\mathbb{R}_{i,0}^{n}$ denote the set of $n$-dimensional vectors $\mu=(\mu_{1},\cdots,\mu_{n})$ satisfying: $i)$ $\mu_{j}\geq 0$ if $j\neq i$, and ii)ii) $\sum_{j\in\mathbb{S}}\mu_{j}=0$. The optimal cost-to-go from time $t$ is defined as $w_{i}(t,x)=\min_{\alpha}J_{i}(t,\alpha)$.
Assumption 1
It is assumed that $1)$ the running cost $c(i,x,\mu)$ is Lipschitz continuous in $x\in\Delta$, with the Lipschitz constant (with respect to $x$) independent of $\mu\in\mathbb{R}_{i,0}^{n}$, $2)$ $c(i,x,\mu)$ is differentiable with respect to $\mu\in\mathbb{R}_{i,0}^{n}$, $3)$ for any $i\in\mathbb{S}$, $x\in\Delta$, the partial derivative $\frac{\partial c}{\partial\mu}(i,x,\mu)$ is Lipschitz continuous with respect to $x\in\Delta$, uniformly in $\mu$, $4)$ $c(i,x,\mu)$ does not depend on $\mu_{i}$, and is uniformly convex: For any $i\in\mathbb{S}$, $x\in\Delta$, $\mu,\mu^{\prime}\in\mathbb{R}_{i,0}^{n}$ with $\mu_{j}\neq\mu_{j}^{\prime}$ for some $j\neq i$, it holds that
$c(i,x,\mu^{\prime})\geq c(i,x,\mu)+\nabla_{\mu}^{\top}c(i,x,\mu)(\mu^{\prime}\!-\!\mu)+\gamma\|\mu^{\prime}\!-\!\mu\|^{2}$
for some positive constant $\gamma$, $5)$ $c(i,x,\mu)$ is superlinear on $\mu_{j},\,j\neq i$: $\lim_{\mu_{j}\to\infty}\frac{c(i,x,\mu)}{\|\mu\|}=\infty$, and $6)$ the terminal cost $\psi(i,x)$ is Lipschitz continuous in $x$. $\square$
The solution to the MFG problem is given by two coupled equations, as outlined in [Diogo_2013]. The first is the backward evolution of $w$, known as the Hamilton-Jacobi (HJ) equation:
@@
The second equation describes the forward evolution of $x(t)$, commonly referred to as the Fokker-Planck (FP) equation:
@@
where $x_{0}=(x_{0,1},\cdots,x_{0,n})\in\Delta$, and
$\alpha_{i}(t)=\textstyle\operatorname*{arg\,min}_{\mu\in\mathbb{R}_{i,0}^{n}}\big(c(i,x(t),\mu)\\ +\textstyle\sum_{j\in\mathbb{S}}\mu_{j}(w_{j}(t,x(t))-w_{i}(t,x(t)))\big),$
representing the optimal state transition. The existence of solutions to (7) and (9) is established in [Diogo_2013, Proposition 4], while their uniqueness is proven in [Diogo_2013, Theorem 2].
The MFG system (7) and (9) may not admit a conventional stationary solution, i.e., one satisfying $\dot{w}(t,x(t))=\dot{x}(t)=0$, for instance, in cases when $H_{i}>0$. Instead, following the approach in [Diogo_2013], we define a stationary solution modulo addition of a constant $\kappa$, characterized by $\dot{w}(t,x(t))=-\kappa\mathbf{1}_{n}$ and $\dot{x}(t)=0$, for which the following conditions hold for all $i\in\mathbb{S}$:
$H_{i}(w(t,x(t)),x(t))$ $=\kappa$ (10a)
$\textstyle\sum_{j\in\mathbb{S}}\alpha_{ji}(t)x_{j}(t)$ $=0.$ (10b)
Note that the existence of such a stationary solution follows from [Diogo_2013, Proposition 5].